OpenStudy (anonymous):

Find the indefinite integral. (Remember to use ln |u| where appropriate.) ∫e^-4x tan(e^-4x)dx i really need help with this...the answer i'm getting is ln|cos e^-4x|

OpenStudy (idku):

\[\Large \int\limits_{ }^{ }~e^{-4x}\tan(e^{-4x})~dx\]

OpenStudy (idku):

Step 1: First set u=-4x (then du = -4 dx, and so you get,) -1/4 du=dx Step 2: set w = e^u dw = e^u du then integrate the tangent....

OpenStudy (idku):

\[\Large \int\limits_{ }^{ }~e^{-4x}\tan(e^{-4x})~dx \]\[u=-4x,~~-\frac{1}{4}du=dx\]\[\Large -\frac{1}{4}\int\limits_{ }^{ }~e^{u}\tan(e^{u})~dx\]

OpenStudy (idku):

\[\Large -\frac{1}{4}\int\limits_{ }^{ }~e^{u}\tan(e^{u})~dx \]\[w=e^u,~~dw=e^u~du\]\[\Large -\frac{1}{4}\int\limits_{ }^{ }\tan(w)~dw\]

OpenStudy (idku):

ohh, it is supposed to be du, where I have e^u, NOT dx..... sorry

OpenStudy (idku):

let me know if you need help integrating the tangent

OpenStudy (anonymous):

the tanget would be cosine right? @idku

OpenStudy (idku):

not exactly.

OpenStudy (idku):

If you don't remember the integral of tangent, then you can do this. \[\int\limits_{ }^{ }\tan(w)dw=\int\limits_{ }^{ }\frac{\sin(w)}{\cos(w)}dw~~~~h=\cos(w),~-dh=\sin(w)\]\[-\int\limits_{ }^{ }1/h~dh=~-\ln(h)+C=-\ln(\cos~w)+C=\ln(\sec~w)+C\]

OpenStudy (idku):

(I brought the exponent inside for the very las last step.

OpenStudy (anonymous):

oh ok i see, so it would be \[\ln|\sec e^-4x|+c\] ?

OpenStudy (idku):

yes, ln ( sec e ^(-4x) ) + C

OpenStudy (idku):

well, or, - ln (cos e^(-4x) ) + C

OpenStudy (anonymous):

ok, thanks

OpenStudy (idku):

anytime